# Portrait3D: 3D Head Generation from Single In-the-wild Portrait Image

### v1

@@ B.1 Align landmark with initialized geometry
In the geometry sculpting stage, landmarks can constrain the face layout in the generation process [13, 22], which effectively solves the multi-face Janus issue. Therefore, we need to obtain landmarks that are aligned with the initialized mesh to guide the geometry generation process.
We first use the face detector (3DDFA [11]) to obtain the 3D landmark of the face from the input image and get seven source keypoints as shown in Figure 10 (b). At the same time, the predicted camera extrinsic parameters are obtained, and this extrinsic component is aligned with the camera parameters in 3D GAN [1]. Next, utilizing the camera position, the imaging plane (blue plane), and the 3D coordinates of the key points on the imaging plane (red points on the blue plane), along with the initialized mesh, we employ the camera position as the origin of the rays (green rays). The directions of the rays are determined by the vectors from the camera position to the keypoints in the imaging plane. We use the Octree-accelerated ray-grid intersection method to compute the intersections of these rays with the initial mesh to obtain the target keypoints, as shown in Figure 10 (a). We then calculate the transformation matrix from the source keypoints to the target points to align the landmarks to the initialized mesh, as shown in Figure 10 (c).
@@ B.2 Texture inpainting details
In the Progressive Texture Inpainting process, we set the camera trajectory to start from the front reference view, symmetrically select the viewpoint for inpainting, and gradually move the camera viewpoint to the back of the head. Specifically, the camera’s horizontal angle moves according to $[0^{\circ},45^{\circ},-45^{\circ},90^{\circ},-90^{\circ},135^{\circ},-135^{\circ},180^{\circ}]$, and the vertical angle is set to $-15^{\circ}$. In this way, we are able to generate a coherent head texture.
@@ B.3 Texture refinement details
Appendix C More ablation
In this section, we show the ablation result of our method in the geometry sculpting stage in Figure 11. (a) Without both ID-aware depth and normal supervision in reference view, the lack of depth results in a flatter head frontal geometry, and a flaky artifact in the geometry, (b) The absence of ID-aware normal map leaves the geometry lacking in detail, (c) Without employing the ID-aware Guidance Enhancement, uneven geometry may arise in the reference viewpoint. This is due to the inconsistency between the geometric supervision under the reference viewpoint and the guidance provided by the SDS. (d) Our full model can generate high-quality geometry compared to (a), (b) and (c).
@@ Appendix C More ablation
Appendix D More result
Figure 12 demonstrates the outstanding performance of our method in generating 3D head from a single in-the-wild portrait image. Our results exhibit high-quality geometry and photorealistic textures. Furthermore, by employing diverse stylized diffusion models, we are capable of generating stylized 3D heads while preserving the portrait’s identity.
Appendix E More comparison
We show more comparison of our method with previous methods in Figure 13. Compared with image-to-3D methods (Magic123 [31], Dreamcraft3D [41], wonder3D [25]), our method has better geometric and texture quality in side view and back view. Also, our method has better facial geometric details compared to Panohead [1].

### v2

@@ Appendix B More implementation details
In the geometry and texture generation stage, we set the resolution of DMTet to 512 to achieve better geometric details. Additionally, the size of the unfolded texture map from the mesh is set to 1024x1024.
@@ B.1 Align landmark with initialized geometry
In the geometry sculpting stage, landmark can constrain the face layout in the generation process (Han et al. 2024; Liu et al. 2023a), which effectively solves the multi-face Janus issue. Therefore, we need to obtain landmarks that are aligned with the initialized mesh to guide the geometry generation process.
We first use face detector (3DDFA (Guo et al. 2020)) to obtain 3D landmark of face from input image and get seven source keypoints as shown in Figure 10 (b). At the same time, the predicted camera extrinsic parameters are obtained, and this extrinsic component is aligned with the camera parameters in 3D GAN (An et al. 2023). Next, utilizing the camera position, the imaging plane (blue plane), and the 3D coordinates of the key points on the imaging plane (red points on the blue plane), along with the initialized mesh, we employ the camera position as the origin of the rays (green rays). The directions of the rays are determined by the vectors from the camera position to the keypoints in the imaging plane. We use the Octree-accelerated ray-grid intersection method to compute the intersections of these rays with the initial mesh to obtain the target keypoints, as shown in Figure 10 (a). We then calculate the transformation matrix from the source keypoints to the target points to align the landmarks to the initialized mesh, as shown in Figure 10 (c).
@@ B.2 Texture inpainting details
In the Progressive Texture Inpainting process, we set the camera trajectory to start from the front reference view, symmetrically select the viewpoint for inpainting, and gradually move the camera viewpoint to the back of the head. Specifically, the camera’s horizontal angle moves according to $[0^{\circ},45^{\circ},-45^{\circ},90^{\circ},-90^{\circ},135^{\circ},-135^{\circ},180^{\circ}]$, and the vertical angle is set to $-15^{\circ}$. In this way, we can generate a coherent head texture.
@@ B.3 Texture refinement details
Appendix C More ablation
The ablation result of our method in the geometry sculpting stage in Figure 11. (a) Without both ID-aware depth and normal supervision in reference view, the lack of depth results in a flatter head frontal geometry, and a flaky artifact in the geometry, (b) The absence of ID-aware normal map leaves the geometry lacking in detail, (c) Without employing the ID-aware Guidance Enhancement, uneven geometry may arise in the reference viewpoint. This is due to the inconsistency between the geometric supervision under the reference viewpoint and the guidance provided by the SDS. (d) Our full model can generate high-quality geometry compared to (a), (b) and (c).
Quantitative ablation result in Table 2 shows that ID-aware Initialization and Guidance greatly enhance the generation quality and identity preservation ability.
Table 2: Quantitative ablations on head generation task. We compute LPIPS under the reference view, ClIP-I and ID score under novel views.
[…]
w/o ID-aware Initialization 0.218 0.5790 0.2296
w/o ID-aware Guidance 0.058 0.6490 0.2984
w/o ID-aware Geo. Supervision 0.050 0.6143 0.4337
w/o Texture Refinement 0.074 0.6368 0.3965
w/o Progressive Tex. Inpainting 0.059 0.6478 0.4269
Full model (ours) 0.039 0.6605 0.4451
Appendix D More result
Figure 13 demonstrates the outstanding performance of our method in generating the 3D head from a single in-the-wild portrait image. Our results exhibit high-quality geometry and photorealistic textures. Furthermore, by employing diverse stylized diffusion models, we are capable of generating stylized 3D heads while preserving the portrait’s identity.
Advantages over Panohead. ID-Sculpt surpasses Panohead in 3 aspects (shown in the Figure below): (1) Better geometry quality (eyes and mouth region). (2) Better texture quality (eye region). (3) Better 3D-consistency: Panohead has weak 3D consistency due to the 2D super-resolution module. We use spatiotemporal textures (Chen, Deng, and Wang 2023) to visualize 3D-consistency, which is obtained by stacking the pixels of a fixed line segment under continuous camera change, where smoothly tilted strips indicate better 3D consistency. Our spatiotemporal textures have fewer noisy patterns.
Figure 12: Comparasion between Panohead and our method.
@@ Appendix D More result
Appendix E More comparison
We show more comparison of our method with previous methods in Figure 14. Compared with image-to-3D methods (Magic123 (Qian et al. 2023), Dreamcraft3D (Sun et al. 2023), wonder3D (Long et al. 2023)), our method has better geometric and texture quality in side view and back view. Also, our method has better facial geometric details compared to Panohead (An et al. 2023).